UT 20130330 Technical Observing - Binocular Testing

Observer: JHill, (Tucson)
Telescope Operator: SAllanson (LBTO)
Additional AO Support: JChristou (LBTO)
Telescope Support: KNewton (LBTO)
Instrument Support: TShih (LBTO)

Summary

This night had scattered thin clouds with good seeing to start out. Both seeing and clouds got worse as the night went on.

We took data to check the DETXY offsets in binocular mode using IRTC.

We took a couple cubes of vibration data with IRTC for Dave A.

We took ~16 stars worth of binocular collimation and pointing data.

The second half was used by LBTI on SX. They got a few things tested, but the majority of the time was lost to clouds and/or a full hard drive on the secondary.

Preparations

Details

Open

01:40 Open - Sky has scattered thin clouds - heavier stuff west of Tucson.

Initial DX Alignment

Fixing the initial coma, and adjusting to a less extreme binodal position before we take collimation data.

02:01 SYNC ACQUIRE Presets to a pointing star at EL=69

02:05 Init DX M2 hexapod as we saw lots of coma. Coma didn't change.

02:10 Adjusting global offsets to remove coma and put smaller RY on DX hexapod, and reduce global astigmatism.

02:21 DX Xglob=-5.00 RYglob=-700

02:22 SYNC ACTIVE Presets

02:25 DX Xglob=-6.00 RYglob=-770 and then coma correction exactly removes that.

Binocular DETXY Offsets

NOTE: AGW2 transform has not been recalibrated since the last remounting of AGW2 on the telescope ~1 week ago

02:40 Preset to BS9134 star guiding off-axis gs=3

02:43 IRS is not running. Why?

[root@tcs2 ~]# service irs restart & [1] 21497 [root@tcs2 ~]# irs: unrecognized service

From the command line: [root@tcs2 ~]# irs Starting IRS server V 3.7

02:49 takepic doesn't give me an IRTC2 image in /newdata/

02:53 irc GetImage works OK - perhaps something in IRAF not updated for IRTC2.

takepic worked fine later at 07:02 UT, so I don't know what was going on here.

02:55 irc GetSeqImages 30

Why does right side have Pointing Origin Base at 0.501,-1.671

That is the default hotspot from PCSInstrument.conf.

03:00 offset 5 0 DETXY ABS which moved -Y on IRTC2 (a known 90 deg mis-orientation of IRTC@RFBG)

lbto@obs2:12 % irc GetImage
irtc2.20130330.030227.fits
----------------------------------------------
lbto@obs2:13 % irc GetSeqImages 30
Resmsg[0]  = IRTC2 GetSeqImages result status: OK
Resmsg[1]  = Number of images acquired: 30 r://irtc2.20130330.00014_cube.fits
Resmsg[2]  = 30 r://irtc2.20130330.00014_cube.fits

03:03 offset -10 0 DETXY ABS which moved +Y on IRTC2

lbto@obs2:14 % irc GetImage
irtc2.20130330.030500.fits
----------------------------------------------
lbto@obs2:15 % irc GetSeqImages 30
Resmsg[0]  = IRTC2 GetSeqImages result status: OK
Resmsg[1]  = Number of images acquired: 30 r://irtc2.20130330.00015_cube.fits
Resmsg[2]  = 30 r://irtc2.20130330.00015_cube.fits

03:03 offset 0 +10 DETXY ABS which moved +X on IRTC2

lbto@obs2:16 % irc GetImage
irtc2.20130330.030839.fits
----------------------------------------------
lbto@obs2:17 % 
lbto@obs2:17 % irc GetSeqImages 30

Note that the pause/resume guiding sequences for these are not clean. Probably because of Issue XXXX that we were investigating 2 nights ago.

03:03 offset 0 -10 DETXY ABS which moved -X on IRTC2

lbto@obs2:18 % irc GetImage
irtc2.20130330.031222.fits
----------------------------------------------
lbto@obs2:19 % irc GetSeqImages 30
Resmsg[0]  = IRTC2 GetSeqImages result status: OK
Resmsg[1]  = Number of images acquired: 30 r://irtc2.20130330.00017_cube.fits
Resmsg[2]  = 30 r://irtc2.20130330.00017_cube.fits

03:10 SX collimation is having trouble image is elongated. It seems to be bouncing in and out of this state. BUT, guider is also reporting 2 FWHM values during this time. ??? INVESTIGATE

03:15 offset 0 0 DETXY ABS

lbto@obs2:20 % irc GetImage
irtc2.20130330.031646.fits
----------------------------------------------
lbto@obs2:21 % irc GetSeqImages 30
Resmsg[0]  = IRTC2 GetSeqImages result status: OK
Resmsg[1]  = Number of images acquired: 30 r://irtc2.20130330.00018_cube.fits
Resmsg[2]  = 30 r://irtc2.20130330.00018_cube.fits

03:15 offset 30 0 DETXY ABS for background sky images

SX Z11 is reasonably stable at 280-210 in 24 minutes - which could be mirror warming. DX Z11 has the same trend plus waves of 200 nm on top with ~10 minute timescale - while we were doing the above offsets.

03:23 offset 0 0 DETXY ABS

03:24 offset 5 0 RADEC ABS -- moved to about 3:30 o'clock

03:25 offset 0 5 RADEC ABS -- moved to about 00:30 o-clock

03:26 offset 0 0 RADEC ABS

03:20 Start Pointing Log - just so we don't forget

03:35 Presets trying to understand Pointing Origin Base - see Night Log -- it is coming from the preset.

03:38 hotspot X=3 Y=3

03:40 hotspot X=-3 Y=-3

03:42 hotspot X=0 Y=0

03:43 hotspot X=-2.0 Y=0

03:44 hotspot X=-1.5 Y=0

High-Speed IRTC data on DX for Ashby

03:49 SYNC ACTIVE preset to mag 2.9 M5 star

03:58 irc GetImage - irtc2.20130330.035816.fits -- full frame

04:00 irc GetImage - irtc2.20130330.040001.fits -- subframe to match the next cube

04:00:44 - 3 ms, 180x180 subframe, H-band, narrow field, guiding
irc GetSeqImages 3000 
Resmsg[0]  = IRTC2 GetSeqImages result status: OK
Resmsg[1]  = Number of images acquired: 3000 r://irtc2.20130330.00020_cube.fits
Resmsg[2]  = 3000 r://irtc2.20130330.00020_cube.fits

04:07 blank sky
 irc GetSeqImages 100
Resmsg[0]  = IRTC2 GetSeqImages result status: OK
Resmsg[1]  = Number of images acquired: 100 r://irtc2.20130330.00021_cube.fits
Resmsg[2]  = 100 r://irtc2.20130330.00021_cube.fits

04:12 - 1ms, 100x100, H-band, medium field, seeing 0.4 arcsec, guiding
lbto@obs2:33 % irc GetImage
irtc2.20130330.041131.fits
----------------------------------------------
lbto@obs2:34 % irc GetSeqImages 3000

04:17 blank sky
irc GetSeqImages 100
Resmsg[0]  = IRTC2 GetSeqImages result status: OK
Resmsg[1]  = Number of images acquired: 100 r://irtc2.20130330.00023_cube.fits
Resmsg[2]  = 100 r://irtc2.20130330.00023_cube.fits

DIMM is bottomed out at 0.6 when the seeing is 0.4. I added a note to Issue 2897.

Single Frame Cube of 3000 Params Cube of 100 Blank
irtc2.20130330.040001.fits irtc2.20130330.00020_cube.fits 3 ms, 180x180 subframe, H-band, narrow field irtc2.20130330.00021_cube.fits
irtc2.20130330.041131.fits irtc2.20130330.00022_cube.fits 1ms, 100x100 subframe, H-band, medium field irtc2.20130330.00023_cube.fits

Binocular Pointing Data

Doug didn't install the script in AOeng or LBTO accounts. Copy the whole /home/dlmiller/idl directory to AOeng. BUT that was very huge! The old AOeng/idl/ is in idl_save. I restored the AOeng account later in the night.

In the mean time we are getting a long guiding and collimation test at EL=68 - I think it crossed the meridian.

/home/aoeng/data/20130330_pointing/

04:45 ptmodel_binoc_collect starting with ACT0363

First star still had -2.5 arcsec in Pointing Model Origin Base. Cleared it by hand on PCSGUI. And then the next preset put it back in.

04:56 ACQUIRE on ACT0397 which put pointing origin base back to 0.501 -1.671 (the repeat of star 3)

Doug's preset used 0.501 -1.671 arcsec (the default value from IRAF).

Steve's IRAF presets all use 0.501 -1.671 unless he explicitly specifies a hotspot. Doug's IDL presets use the last hotspot numbers that Steve sent with IRAF. The answer is that Steve/IRAF and John/IDL were using the same proxies, and Doug doesn't say "Clear Hotspot" when he sends a preset. I send an email to Doug.

aoeng@obs2:26 % irc -show
Proxy name: irc

IIF connections:
irc : IRTC at bentGregorianFront left
LBTI LBTO-INDI Gateway : LBTI at bentGregorianCenter both
IRC : IRTC at bentGregorianFront left
DIMM : DIMM at prime right
IRCR : IRTC at bentGregorianFront right
IRCL : LUCIFER at bentGregorianFront left
PEPSI_PFU_left : PEPSIFIBER at bentGregorianRearFiberFeed left
PEPSI_PFU_right : PEPSIFIBER at bentGregorianRearFiberFeed right
PEPSI_PFU_both : PEPSIFIBER at bentGregorianRearFiberFeed both
PEPSI_POL_left : PEPSI at directGregorian left
PEPSI_POL_right : PEPSI at directGregorian right
PEPSI_POL_both : PEPSI at directGregorian both
LBTIDD : LBTI at bentGregorianCenter both

Camera assignments:
IRC1 : IRTC1
IRC2 : IRTC2
IRTC1GUI : IRTC1
IRTC2GUI : IRTC2

05:11 ACT0284 was a close binary star.

Can huge astigmatism drive the spots into the edge of the field stop and make high-order zernikes?

05:17 John's PCSGUI core dumped as telescope. Added a note to IT 3332

05:18 see the ghost in the acquisition image at this time. The ghost image shows that the edge of the shell is wrinkly.

I have submitted IT 4585 about the wrinkly pupil.

05:20 ACQUIRE presets to correct pointing

05:21 Rest and re-flatten the right shell.

05:24 ptmodel_collect, start=6

05:43 ptmodel_binoc_collect, start=11 which is ACT0245

IDL> ptmodel_binoc_collect, start=11
  
      # Target star: 0   ACT0245
        14:45:21.548 +10:35:56.39 RADEC 2000.0 J2000 -16.0 6.0 6.26 R 1.20 B_V 1.65
      # Guide star: 1
        14:45:36.494 +10:36:05.25 RADEC 2000.0 J2000 8.0 -6.0 14.58 R 1.09 B_V 0.63
  
      preset, ACT0245, MODE=ACTIVE, GS=1, PA=87.7297     (iif_preset)
  
  
      # Target star: 0   ACT0245
        14:45:21.548 +10:35:56.39 RADEC 2000.0 J2000 -16.0 6.0 6.26 R 1.20 B_V 1.65
      # Guide star: 1
        14:45:36.494 +10:36:05.25 RADEC 2000.0 J2000 8.0 -6.0 14.58 R 1.09 B_V 0.63
  
      preset, ACT0245, MODE=ACTIVE, GS=1, PA=87.7297     (iif_preset)
  
           232 nm rms error in image left_wfscimage000235.fits
   => Collimated in 1 iteration
===> Active Optics loop is not running

05:46 ptmodel_binoc_collect, start=11 which is ACT0245

06:12 ACT0118 is through cloud --- the pointing is OK, but maybe wfs won't converge.

      # Target star: 0   ACT0118
        11:55:56.977 -24:55:32.80 RADEC 2000.0 J2000 -8.0 -46.0 7.07 R 1.16 B_V 1.65
      # Guide star: 5
        11:56:16.074 -24:53:42.02 RADEC 2000.0 J2000 0.0 0.0 14.63 R 1.68 B_V 0.63
  
      preset, ACT0118, MODE=ACTIVE, GS=5, PA=66.8179     (iif_preset)
  
           617 nm rms error in image left_wfscimage000257.fits
===> WARNING in recon_create_recon: Source Extractor found too few S-H spots: 1
===> WARNING in recon_determine_zernikes: wavefront not reconstructed 
         => Could not process /Repository/GCSfiles/dailyimages//left_wfscimage000257.fits
              Probably could not find info file
              due to bad WFSC image
  
   => Collimated in 1 iteration
% Program caused arithmetic error: Floating illegal operand

First two stars in the table have suspect pointing origin base values. First one is certainly off, but second one could be OK.

num Star Name GStar WFSC Image WFSC Image
1 ACT0363 2 left_wfscimage000193.fits right_wfscimage000172.fits
2 WT10_406 3 left_wfscimage000196.fits right_wfscimage000176.fits
num Star Name GStar WFSC Image WFSC Image
3 ACT0397 1 left_wfscimage000202.fits right_wfscimage000179.fits
4 WT10_348 1 left_wfscimage000206.fits right_wfscimage000182.fits
5 ACT0284 6 left_wfscimage000212.fits right_wfscimage000188.fits
num Star Name GStar WFSC Image WFSC Image
6 ACT0218 6 left_wfscimage000224.fits right_wfscimage000193.fits
7 WT10_116 3 left_wfscimage000229.fits right_wfscimage000198.fits
8 WT10_116 3 left_wfscimage000230.fits right_wfscimage000200.fits
9 ACT0188 3 left_wfscimage000230.fits right_wfscimage000200.fits
num Star Name GStar WFSC Image WFSC Image
11 ACT0245 1 left_wfscimage000238.fits right_wfscimage000203.fits
12 WT10_341 4 left_wfscimage000240.fits right_wfscimage000206.fits
13 WT10_326 1 left_wfscimage000244.fits right_wfscimage000209.fits
14 WT10_296 1 left_wfscimage000248.fits right_wfscimage000214.fits
15 ACT0262 1 left_wfscimage000251.fits right_wfscimage000217.fits
16 WT10_292 1 left_wfscimage000251.fits right_wfscimage000217.fits
17 WT10_205 2 left_wfscimage000255.fits right_wfscimage000221.fits
18 ACT0118 5 left_wfscimage000257.fits right_wfscimage000224.fits

LBTI AO testing on SX

06:15 Handover to LBTI

LBTI is now sending the correct proper motions.

2013-03-30T06:18:35.297139+00:00 tcs2 LBT_PCS: PCS. setEquatorialTarget - MidpointX: -100 MidpointY: -100
2013-03-30T06:18:35.297158+00:00 tcs2 LBT_PCS: PCS. setEquatorialTarget - SXSeparation: 0 DXSeparation: 0
2013-03-30T06:18:35.338317+00:00 tcs2 LBT_PCS: setEquatorialTarget. 172.5623310098 left 15:01:55.165958495 +40:23:21.751604568 Mode:PARALLACTIC Angle:0 Wave: 0.8000000119209
2013-03-30T06:18:35.338350+00:00 tcs2 LBT_PCS: Proper Motion RA(rad): -1.946469521605e-07 Proper Motion Dec(rad): -1.399131018519e-07
2013-03-30T06:18:35.338364+00:00 tcs2 LBT_PCS: PM CorrRA: 15:01:55.130514273 PM CorrDec: +40:23:21.369441648
2013-03-30T06:18:35.338648+00:00 tcs2 LBT_PCS: setEquatorialTarget.  mcspuFS: LCBG rotIndex: -10
2013-03-30T06:18:35.359962+00:00 tcs2 LBT_PCS: setEquatorialTarget. 172.5624229274 right 15:01:55.165958495 +40:23:21.751604568 Mode:PARALLACTIC Angle:0 Wave: 0.8000000119209
2013-03-30T06:18:35.359993+00:00 tcs2 LBT_PCS: Proper Motion RA(rad): -1.946469521605e-07 Proper Motion Dec(rad): -1.399131018519e-07
2013-03-30T06:18:35.360019+00:00 tcs2 LBT_PCS: PM CorrRA: 15:01:55.130514273 PM CorrDec: +40:23:21.369441648
2013-03-30T06:18:35.360329+00:00 tcs2 LBT_PCS: setEquatorialTarget.  mcspuFS: RCBG rotIndex: -10
2013-03-30T06:18:35.369218+00:00 tcs2 LBT_PCS: TrajectoryGenerator.   Clear polys and RotIndex. Side: 0 rotIndexSX: -10 rotIndexDX: -10
2013-03-30T06:18:35.369260+00:00 tcs2 LBT_PCS: EncoderFeedback.   Clear polys and RotIndex. Side: 0 rotIndexSX: -10 rotIndexDX: -10

06:25 Stop pointing log

07:30 SX Shell RIPS and won't recover.

07:50 Authorize as IRTC@RFBG, but can't find an M5 star.

08:00 Close for clouds

08:01 Authorize as LBTI on left so Vidhya can test software commands in closed dome.

-- JohnHill - 30 Mar 2013
Topic revision: r4 - 02 Apr 2013, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback